OpenStudy (anonymous):

use the definition of derivative to find the derivative of f(x)=1/x+3

OpenStudy (anonymous):

x^-1+3=(-1)x^-2=-1/x^2

OpenStudy (anonymous):

what is the formula for the definition of derivatives?

OpenStudy (nikvist):

\[f'(x)=\lim_{h\rightarrow0}\frac{f(x+h)-f(x)}{h}\]